Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
SASE1 Status and SASE3 Readiness Technical integration and coordination
  Joint Operation and Readiness Logbook, Contact: T.Haas, A. Violante, G. Wellenreuther  Not logged in ELOG logo
Message ID: 581     Entry time: 26 Nov 2018, 10:35
Author: Adriano Violante 
Location: General 
Severity: Information 
Subject: Status report after dispatch meeting 09.30 

SPB

  • KB system commissioning ongoing
  • Agipd commissioning stopped
  • Radiation tests preparation ongoing.

FXE

SQS

SCS

MID

  • TueV Test successful.
  • Mono commissioning issue: the pico motors position is not reproducible. This is a show stopper for the Monochromator commissionig. Not clear whether also FXE has the same kind of problem. AE has been informed and will try to support taking into account also the requests by SASE3 for this week (commissionig for the users run)
  • AGIPD commissioning is ongoing. Chiller has to be replaced due to a leak.

HED

  • TueV Test for the optics hutch successful
  • Cabling ongoing. 
  • Fw export and crates/loops preparation ongoing
  • Screen is needed for first beam.

XRO

XPD

VAC

AE

  • OCD: intervention request during the WE to fix a train id issue. Should be ok now but the problem still needs to be reproduced and undertstood.

ITDM

  • Network intervention needs to be planned for the beginning of January. KW is clarifying the details. Will be included in the shutdown planning.

CAS

  • Release tests ongoing. Small issue with the DAQ has been found. Fix already in place. Final tst of the fix is being performed. 

AoB

ELOG V3.1.4-7c3fd00